A boat cruises 36 miles against a 3mph current and 36 miles back against the same current. The round trip takes 5 hours. What is the speed of the boat in still water?
Answer by mananth(16949)

let speed in still water = xmph
current speed = 3mph
upstream speed = x-3 mph
downstream speed = x+3 mph
...
time = distance / speed
36/(x+3)+36/(x-3)=5
36(x-3)+36(x+3)/(x+3)(x-3)=5
36x-108+36x+108=5(x^2-9)
36x+36x=5x^2-45
5x^2-72x-45=0
5x^2-75x+3x-45=0
5x(x-15)+3(x-15)=0
(x-15)(5x+3)=0
x=15 OR x=-3/5 ignore
x= 15 boat speed in still water
,...
m.ananth@hotmail.ca